OpenStudy (precal):

Find all functions f whose derivatives is f prime = x+1

OpenStudy (precal):

\[f \prime=x+1\]

OpenStudy (bahrom7893):

Integrate x+1: f(x) = x^2/2 + x + C

OpenStudy (bahrom7893):

 C is any constant